consider an infinitely long annular cylinder of inner radius ri= 290 mm and outer radius ro= 435 mm. the cylinder is subjected to an internal pressure p= 100 mpa. the outer wall of the cylinder is constrained against any movement. the young's modulus of the cylinder material is e= 192 gpa and the poisson's ratio is ν= 0.3. determine the magnitude of the radial pressure (in mpa) at the outer radius
1
Expert's answer
2021-10-28T08:53:10-0400

"E_1=\mu E_2,"

"2\pi r_1(p_1+E_1)=2\pi r_2(p_2+E_2),"

"p_2=(p_1+E_1)\frac{r_1}{r_2}-E_2,"

"p_2=87~MPa."